(3) Given (r y 1)dr (y--5)dy 0. Find constants h and k so that the substitutions r u h and y + k result in a homogencous differential equation. Now solve the differential equation.
Answer #1

③ Sol utioGiven differential equation is and dx - du, dy dv dv dy dx. du 4 Equaty.on ④ becomes homogeneous if h+k-1 0 By solv

Now, Eqn⑤→ due utv Dividing Nr. and Dr. by u,we get du el v dn 2 dd 2 Taking variable and seperable

Tanta te 2. すな).dz = Rolf. (z )kl but u = x-h , v = y-k [.. From(2) ] General solution for the given differential equation is